# Overcoming the Shadow: Bending Airy Beams for Radiative Near-Field Multi-User Access in Half-Space Blockage Scenarios

### v1

@@ Abstract
The move to next-generation wireless communications with extremely large-scale antenna arrays (ELAAs) brings the communications into the radiative near-field (RNF) region, where distance-aware focusing is feasible. However, high-frequency RNF links are highly vulnerable to blockage in indoor environments dominated by half-space obstacles (walls, corners) that create knife-edge shadows. Conventional near-field focused beams offer high gain in line-of-sight (LoS) scenarios but suffer from severe energy truncation and effective-rank collapse in shadowed regions, making hardware remedies such as reconfigurable intelligent surfaces (RIS) impractical. We propose a beamforming strategy that exploits the auto-bending property of Airy beams to mitigate half-space blockage without additional hardware. The Airy beam is designed to “ride” the diffraction edge, accelerating its main lobe into the shadow to restore connectivity. Our contributions are threefold: (i) a Green’s function-based RNF multi-user channel model that analytically reveals singular-value collapse behind knife-edge obstacles; (ii) an Airy analog beamforming scheme that optimizes the bending trajectory to recover the effective channel rank; and (iii) an Airy null-steering method that aligns oscillatory nulls with bright-region users to suppress interference in mixed shadow/bright scenarios. Simulations show that the proposed edge-riding Airy strategy achieves an SNR improvement of over 20 dB and restores full-rank connectivity in shadowed links compared to conventional RNF focusing, virtually eliminating outage in geometric shadows and increasing multi-user spectral efficiency by approximately 35% under typical indoor ELAA configurations. These results demonstrate robust RNF multi-user access in half-space blockage scenarios without relying on RIS.
@@ I Introduction
Recent years have witnessed a surge of interest in Non-Diffracting (ND) beams for this purpose. The seminal work by Reddy and Jornet et al. [16] demonstrated the potential of THz Bessel beams to “self-heal” after encountering partial blockage, triggering a wave of research into structured light for wireless communications. Subsequent studies have expanded this frontier: Miao et al. [19] realized extended non-diffractive ranges using dielectric metalenses, while Yang et al. [26] explored flexible antenna arrays to dynamically shape radiation patterns. In our recent work [18], we provided a quantitative roadmap for the engineering of linear ND beams (specifically Bessel and Mathieu beams) using planar phased arrays, establishing the fundamental link-level gain regimes for these structured wavefronts.
However, a critical limitation remains: Bessel and Mathieu beams, while robust to partial occlusion, still propagate along linear trajectories. They cannot reach deep into the geometric shadow of large, optically opaque obstacles. This limitation has catalyzed the exploration of Airy beams, which possess a unique auto-bending property capable of curving around obstacles. Guerboukha et al. [17] provided the first experimental demonstration of curving THz data links, establishing a high ground for this emerging field. Subsequently, the potential of Airy beams has been rapidly explored across various domains: Darsena et al. [22] analyzed their fundamental performance limits in near-field channels; Lee et al. [21] experimentally validated high-speed transmission using 3D-printed lenses; and Zhang et al. [24] proposed convex optimization frameworks for bending beam generation. Beyond standard communications, the unique wavefronts of Airy beams have also been exploited for physical layer security via wavefront hopping [20] and high-resolution phase imaging through scattering media [23, 27].
@@ III-A Simulation Geometry and Parameters
  • UE1 (Target User): Fixed at coordinates $\mathbf{p}_{1}=(-5\lambda,250\lambda)$. Remark: While this location is Line-of-Sight (LoS) in this baseline section, it is strategically chosen to fall within the “geometric shadow” region ($x<0$) once the half-space obstacle is introduced in Section IV.
  • UE2 (Interfering User): Located at a different depth $z_{2}=300\lambda$. This user scans laterally along the $x$-axis from $x_{2}=-15\lambda$ (deep shadow side) to $10\lambda$ (bright side), crossing the angular line-of-sight of UE1.
@@
Figure 2: Baseline performance of the RNF-MU system with conventional focusing in free space. (a) System Geometry: UE1 fixed at $(-5\lambda,250\lambda)$ and UE2 scanning at $z=300\lambda$. (b) Condition Number $\kappa(\mathbf{H}_{\text{eff}})$: A sharp singularity ($\kappa_{\max}\approx 158$) occurs at $x_{2}=-6\lambda$, corresponding to angular alignment with UE1. (c) Common SINR: Equalized SINR profile showing degradation at the angular collision. (d) Sum-Rate: Minimum throughput observed at the singularity point. (e)–(f) Angular Analysis: SINR and $\kappa$ plotted against UE2’s angle $\theta_{2}$, confirming the overlap occurs at $\theta_{2}\approx\theta_{1}\approx-1.15^{\circ}$. (g)–(i) Field Intensity Distributions: Normalized field power $|E(x,z)|^{2}$ for UE2 located at $x_{2}=-10\lambda$, $0\lambda$, and $10\lambda$, respectively, visualizing the distinct focal spots.
@@ III-C1 Condition Number and Angular Singularity
Fig. 2(b) plots the condition number $\kappa(\mathbf{H}_{\text{eff}})$ against the lateral position of UE2. A critical observation is that the peak condition number ($\kappa\approx 158$) occurs at $x_{2}=-6\lambda$, not at $x_{2}=-5\lambda$. This shift is physically significant:
  • UE1 is at $(-5\lambda,250\lambda)$, corresponding to an angle $\theta_{1}\approx\text{atan}(-5/250)\approx-1.15^{\circ}$.
  • The singularity occurs when UE2 (at $z_{2}=300\lambda$) aligns with this angle: $x_{2}=300\lambda\cdot\tan(\theta_{1})=-6\lambda$.
Thus the collision occurs when UE2 shares the same ray angle with UE1, not when $x_{2}$ equals $x_{1}$. This indicates that for the given aperture ($D\approx 31\lambda$) and depth separation ($50\lambda$), traditional focusing relies primarily on angular separation to maintain orthogonality. The longitudinal resolution is insufficient to decorrelate users located on the same ray, leading to a “co-linear singularity.”
@@ III-C2 Spectral Efficiency
IV Half-Space Blockage: Double-Shadow Users and Airy-All Analog Beamforming
@@ IV-A Half-Space Blockage Scenario and Singular Value Collapse
  • UE1 (Target): Remains fixed at $(x_{1},z_{1})=(-5\lambda,250\lambda)$, consistent with the baseline geometry (except now obstructed).
  • UE2 (Scanning): Is located at depth $z_{2}=300\lambda$ but is now scanned laterally from $x_{2}=-15\lambda$ (deep shadow) to $x_{2}=-1\lambda$ (near the diffraction edge). This range is specifically chosen to investigate performance deep inside the shadow where traditional links fail.
@@
IV-B Proposed Airy-All Analog Beamforming Strategy
To overcome the shadow, we propose an “Airy-All” analog beamforming strategy. Unlike traditional beams that focus energy on a straight line, Airy beams follow a curved parabolic trajectory, enabling them to “ride” the diffraction edge and accelerate into the shadow zone.
@@ IV-B1 Fixed Geometry-Based Codebook Design
Figure 3: Performance comparison in the double-shadow scenario ($x_{1}=-5\lambda$, fixed; $x_{2}$ scanning inside shadow). (a) System geometry showing the knife-edge blockage at $z=150\lambda$. (b) Sum-rate performance: Airy beams (red) maintain a robust rate floor in the deep shadow where traditional beams (blue) suffer outage. (c) Sum-rate gain of Airy over traditional focusing, showing gains up to 18 bps/Hz. (d) Common SINR: Airy beams provide a usable SINR floor despite interference, whereas traditional SINR collapses. (e) Condition number: Note that while Airy has a higher $\kappa$ near collision, it provides non-zero singular values, enabling communication. (f) Steering angles used for the Airy beams, following simple geometric logic.
@@ IV-C1 SINR Resilience in Deep Shadow
Fig. 3(d) compares the Common SINR of the Traditional and Airy-All schemes.
@@ IV-C2 Mechanism of Recovery: Active Bending
  • Deep Shadow Recovery (Top Row): Fig. 4(a) and (b) correspond to a large separation ($x_{2}=-11\lambda$). Traditional beams (a) are sharply truncated, creating a dark triangular shadow zone. In contrast, Airy beams (b) exhibit a clear curved trajectory. The main lobe accelerates transversely, effectively “turning the corner” of the obstacle to illuminate the deep shadow user UE2.
  • Recovery under Angular Collision (Bottom Row): Fig. 4(c) and (d) depict the scenario where UE2 is close to UE1 ($x_{2}\approx-5\lambda$). Here, both users are shadowed. While traditional beams (c) result in a complete double outage, the Airy beams (d) manage to illuminate both locations. Although the beams spatially overlap near the obstacle edge, they retain sufficient distinctness in the depth domain to allow the ZF precoder to separate the streams.
@@ IV-C4 Summary
In summary, the proposed Airy-All strategy, even with a fixed and sub-optimal parameter configuration, successfully mitigates half-space blockage. By exploiting the auto-bending property, it transforms the “hard” geometric shadow into a “soft” coverage zone, providing a reliable fallback for RNF multi-user communications.
@@ V-A Scenario: The Interference Challenge
  • UE1 (Shadow): Located at $(-5\lambda,250\lambda)$, served by an Airy beam.
  • UE2 (Bright - Hard Case): Located at $(+3.5\lambda,300\lambda)$. This position is termed a “hard case” because it lies close to the diffraction boundary, where the oscillating tail of UE1’s Airy beam is strong.
  • Precoding Strategy: UE2 uses a fixed traditional beam (optimal for LoS). The optimization focuses solely on UE1’s Airy beam parameters.
If we apply the Airy-Geo strategy (geometric steering) from Section IV, the main lobe serves UE1, but the side-lobes heavily pollute UE2’s channel. This degrades the condition number of $\mathbf{H}_{\text{eff}}$, causing the ZF precoder to amplify noise ($N_{0}=10^{-3}$) and limiting the sum-rate.
@@ V-C Proposed Solution: Coarse-to-Fine Search
Output: Optimal UE1 Airy parameters $(B^{\star},F^{\star},\theta_{1}^{\star})$.
@@
4 Fix UE2 beam (traditional) and compute its channel column $\mathbf{h}_{2}$
@@ 7 $R_{\text{best}}\leftarrow$-$\infty$
11        Generate candidate UE1 Airy beam $E_{1}(x)$ with parameters $(B,F,\theta_{\text{geo}}+\Delta\theta)$
@@ 19           $R_{\text{best}}\leftarrow R$
34        Generate candidate UE1 Airy beam $E_{1}(x)$ with parameters $(B,F,\theta_{\text{best}}+\Delta\theta)$
@@ V-D Complexity Analysis
Let $N_{\text{coarse}}=|\mathcal{B}_{c}||\mathcal{F}_{c}||\Delta\Theta_{c}|$ and $N_{\text{fine}}=|\mathcal{B}_{f}||\mathcal{F}_{f}||\Delta\Theta_{f}|$ denote the number of candidate parameter tuples evaluated in the coarse and fine search stages, respectively, consistent with the sets defined in Algorithm 1. For each candidate tuple, the dominant computational cost lies in the evaluation of the Fresnel propagation operator $\mathcal{P}_{z}$ (see Eq. (3)). Implementing this operator via the FFT-based Angular Spectrum Method on a transverse aperture grid of size $N_{x}$ (where $N_{x}\geq N$ to ensure sufficient sampling resolution) yields a propagation complexity of $\mathcal{O}(N_{x}\log N_{x})$. Consequently, the total computational complexity of the proposed algorithm scales as:
@@
It is worth noting that since the traditional focused beam for the Line-of-Sight (LoS) user (UE2) is fixed during the optimization of the shadowed user (UE1), its effective channel column $\mathbf{h}_{2}$ can be precomputed once. This reduces the per-iteration cost solely to the update of the Airy beam column $\mathbf{h}_{1}$.
@@
Figure 5: Airy-Null optimization results in the mixed Shadow-Bright scenario. (a) Geometry showing UE1 in shadow and UE2 in bright region. (b) Sum-rate vs. steering angle $\Delta\theta$: The peak occurs at $\Delta\theta\approx-2.8^{\circ}$, not at the geometric zero. (c) Condition number $\kappa$ vs. $\theta$: A sharp dip in $\kappa$ coincides with the sum-rate peak, indicating improved orthogonality. (d)-(e) Comparison of Sum-Rate and Condition Number for Trad-All, Airy-Geo, and Airy-Opt strategies. (f) Analog coupling strengths, highlighting the interference suppression capability of the optimized design.

### v2

# Airy Beamforming for Radiative Near-Field MU-XL-MIMO: Overcoming Half-Space Blockage

@@ Abstract
The move to next-generation wireless communications with extremely large-scale antenna arrays (ELAAs) brings the communications into the radiative near-field (RNF) region, where distance-aware focusing is feasible. However, high-frequency RNF links are highly vulnerable to blockage in indoor environments dominated by half-space obstacles (walls, corners) that create knife-edge shadows. Conventional near-field focused beams offer high gain in line-of-sight (LoS) scenarios but suffer from severe energy truncation and effective-rank collapse in shadowed regions, often necessitating the deployment of auxiliary hardware such as Reconfigurable Intelligent Surfaces (RIS) to restore connectivity. We propose a beamforming strategy that exploits the auto-bending property of Airy beams to mitigate half-space blockage without additional hardware. The Airy beam is designed to “ride” the diffraction edge, accelerating its main lobe into the shadow to restore connectivity. Our contributions are threefold: (i) a Green’s function-based RNF multi-user channel model that analytically reveals singular-value collapse behind knife-edge obstacles; (ii) an Airy analog beamforming scheme that optimizes the bending trajectory to recover the effective channel rank; and (iii) an Airy null-steering method that aligns oscillatory nulls with bright-region users to suppress interference in mixed shadow/bright scenarios. Simulations show that the proposed edge-riding Airy strategy achieves a Signal-to-Noise Ratio (SNR) improvement of over 20 dB and restores full-rank connectivity in shadowed links compared to conventional RNF focusing, virtually eliminating outage in geometric shadows and increasing multi-user spectral efficiency by approximately 35% under typical indoor ELAA configurations. These results demonstrate robust RNF multi-user access in half-space blockage scenarios without relying on RIS.
@@ I Introduction
Recent years have witnessed a surge of interest in Non-Diffracting (ND) beams for this purpose. The seminal work by Reddy and Jornet et al. [16] demonstrated the potential of THz Bessel beams to “self-heal” after encountering partial blockage, triggering a wave of research into structured light for wireless communications. Subsequent studies have expanded this frontier: Miao et al. [19] realized extended non-diffractive ranges using dielectric metalenses, while Yang et al. [26] explored flexible antenna arrays to dynamically shape radiation patterns. In our recent work [18], we provided a quantitative roadmap for the engineering of linear ND beams (specifically Bessel and Mathieu beams) using planar phased arrays, establishing the fundamental link-level gain regimes for these structured wavefronts [18]. Complementing these theoretical studies, Bodet et al. [30] provided essential sub-terahertz near-field channel measurements, verifying the propagation characteristics of diffractive beams.
However, a critical limitation remains: Bessel and Mathieu beams, while robust to partial occlusion, still propagate along linear trajectories. They cannot reach deep into the geometric shadow of large, optically opaque obstacles. Subsequently, the potential of Airy beams has been rapidly explored across various domains: Darsena et al. [22] analyzed their fundamental performance limits in near-field channels; Lee et al. [21] experimentally validated high-speed transmission using 3D-printed lenses; and Zhang et al. [24] proposed convex optimization frameworks for bending beam generation. In specific application scenarios, Zhao et al. [25] evaluated the efficacy of Airy beams for Terahertz wireless data centers, while Hu et al. [29] demonstrated robust underwater optical communications using RGB Airy transmitters. Beyond standard communications, the unique wavefronts of Airy beams have also been exploited for physical layer security via wavefront hopping [20], complex spatiotemporal wavepacket dynamics [28], and high-resolution phase imaging through scattering media [23, 27].
@@ III-A Simulation Geometry and Parameters
  • UE-1 (Target User): Fixed at coordinates $\mathbf{p}_{1}=(-5\lambda,250\lambda)$. Remark: While this location is Line-of-Sight (LoS) in this baseline section, it is strategically chosen to fall within the “geometric shadow” region ($x<0$) once the half-space obstacle is introduced in Section IV.
  • UE-2 (Interfering User): Located at a different depth $z_{2}=300\lambda$. This user scans laterally along the $x$-axis from $x_{2}=-15\lambda$ (deep shadow side) to $10\lambda$ (bright side), crossing the angular line-of-sight of UE-1.
@@
Figure 2: Baseline performance of the RNF-MU system with conventional focusing in free space. (a) System Geometry: UE-1 fixed at $(-5\lambda,250\lambda)$ and UE-2 scanning at $z=300\lambda$. (b) Condition Number $\kappa(\mathbf{H}_{\text{eff}})$: A sharp singularity ($\kappa_{\max}\approx 158$) occurs at $x_{2}=-6\lambda$, corresponding to angular alignment with UE-1. (c) Common SINR: Equalized SINR profile showing degradation at the angular collision. (d) Sum-Rate: Minimum throughput observed at the singularity point. (e)–(f) Angular Analysis: SINR and $\kappa$ plotted against UE-2’s angle $\theta_{2}$, confirming the overlap occurs at $\theta_{2}\approx\theta_{1}\approx-1.15^{\circ}$. (g)–(i) Field Intensity Distributions: Normalized field power $|E(x,z)|^{2}$ for UE-2 located at $x_{2}=-10\lambda$, $0\lambda$, and $10\lambda$, respectively, visualizing the distinct focal spots.
@@ III-C1 Condition Number and Angular Singularity
Fig. 2(b) plots the condition number $\kappa(\mathbf{H}_{\text{eff}})$ against the lateral position of UE-2. A critical observation is that the peak condition number ($\kappa\approx 158$) occurs at $x_{2}=-6\lambda$, not at $x_{2}=-5\lambda$. This shift is physically significant:
  • UE-1 is at $(-5\lambda,250\lambda)$, corresponding to an angle $\theta_{1}\approx\text{atan}(-5/250)\approx-1.15^{\circ}$.
  • The singularity occurs when UE-2 (at $z_{2}=300\lambda$) aligns with this angle: $x_{2}=300\lambda\cdot\tan(\theta_{1})=-6\lambda$.
Thus the collision occurs when UE-2 shares the same ray angle with UE-1, not when $x_{2}$ equals $x_{1}$. This indicates that for the given aperture ($D\approx 31\lambda$) and depth separation ($50\lambda$), traditional focusing relies primarily on angular separation to maintain orthogonality. The longitudinal resolution is insufficient to decorrelate users located on the same ray, leading to a “co-linear singularity.”
@@ III-C2 Spectral Efficiency
IV Half-Space Blockage: Double-Shadow Users and Airy-Geo Analog Beamforming
@@ IV-A Half-Space Blockage Scenario and Singular Value Collapse
  • UE-1 (Target): Remains fixed at $(x_{1},z_{1})=(-5\lambda,250\lambda)$, consistent with the baseline geometry (except now obstructed).
  • UE-2 (Scanning): Is located at depth $z_{2}=300\lambda$ but is now scanned laterally from $x_{2}=-15\lambda$ (deep shadow) to $x_{2}=-1\lambda$ (near the diffraction edge). This range is specifically chosen to investigate performance deep inside the shadow where traditional links fail.
@@
IV-B Proposed Airy-Geo Analog Beamforming Strategy
To overcome the shadow, we propose an “Airy-Geo” analog beamforming strategy. Unlike traditional beams that focus energy on a straight line, Airy beams follow a curved parabolic trajectory, enabling them to “ride” the diffraction edge and accelerate into the shadow zone.
@@ IV-B1 Fixed Geometry-Based Codebook Design
Figure 3: Performance comparison in the double-shadow scenario ($x_{1}=-5\lambda$, fixed; $x_{2}$ scanning inside shadow). (a) System geometry showing the knife-edge blockage at $z=150\lambda$. (b) Sum-rate performance: Airy beams (red) maintain a robust rate floor in the deep shadow where traditional beams (blue) suffer outage. (c) Sum-rate gain of Airy over traditional focusing, showing gains up to 4 bps/Hz. (d) Common SINR: Airy beams provide a usable SINR floor despite interference, whereas traditional SINR collapses. (e) Condition number: Note that while Airy has a higher $\kappa$ near collision, it provides non-zero singular values, enabling communication. (f) Steering angles used for the Airy beams, following simple geometric logic.
@@ IV-C1 SINR Resilience in Deep Shadow
Fig. 3(d) compares the Common SINR of the Traditional and Airy-Geo schemes.
@@ IV-C2 Mechanism of Recovery: Active Bending
  • Deep Shadow Recovery (Top Row): Fig. 4(a) and (b) correspond to a large separation ($x_{2}=-11\lambda$). Traditional beams (a) are sharply truncated, creating a dark triangular shadow zone. In contrast, Airy beams (b) exhibit a clear curved trajectory. The main lobe accelerates transversely, effectively “turning the corner” of the obstacle to illuminate the deep shadow user UE-2.
  • Recovery under Angular Collision (Bottom Row): Fig. 4(c) and (d) depict the scenario where UE-2 is close to UE-1 ($x_{2}\approx-5\lambda$). Here, both users are shadowed. While traditional beams (c) result in a complete double outage, the Airy beams (d) manage to illuminate both locations. Although the beams spatially overlap near the obstacle edge, they retain sufficient distinctness in the depth domain to allow the ZF precoder to separate the streams.
@@ IV-C4 Summary
In summary, the proposed Airy-Geo strategy, even with a fixed and sub-optimal parameter configuration, successfully mitigates half-space blockage. By exploiting the auto-bending property, it transforms the “hard” geometric shadow into a “soft” coverage zone, providing a reliable fallback for RNF multi-user communications.
@@ V-A Scenario: The Interference Challenge
  • UE-1 (Shadow): Located at $(-5\lambda,250\lambda)$, served by an Airy beam.
  • UE-2 (Bright - Hard Case): Located at $(+3.5\lambda,300\lambda)$. This position is termed a “hard case” because it lies close to the diffraction boundary, where the oscillating tail of UE-1’s Airy beam is strong.
  • Precoding Strategy: UE-2 uses a fixed traditional beam (optimal for LoS). The optimization focuses solely on UE-1’s Airy beam parameters.
If we apply the Airy-Geo strategy (geometric steering) from Section IV, the main lobe serves UE-1, but the side-lobes heavily pollute UE-2’s channel. This degrades the condition number of $\mathbf{H}_{\text{eff}}$, causing the ZF precoder to amplify noise ($N_{0}=10^{-3}$) and limiting the sum-rate.
@@ V-B Problem Formulation: Joint Parameter Optimization
Output: Optimal UE-1 Airy parameters $(B^{\star},F^{\star},\theta_{1}^{\star})$.
@@
4 Fix UE-2 beam (traditional) and compute its channel column $\mathbf{h}_{2}$
@@ 7 $R_{\text{best}}\leftarrow$-$\infty$
11        Generate candidate UE-1 Airy beam $E_{1}(x)$ with parameters $(B,F,\theta_{\text{geo}}+\Delta\theta)$
@@ 19           $R_{\text{best}}\leftarrow R$
34        Generate candidate UE-1 Airy beam $E_{1}(x)$ with parameters $(B,F,\theta_{\text{best}}+\Delta\theta)$
@@ V-D Complexity Analysis
Let $N_{\text{coarse}}=|\mathcal{B}_{c}||\mathcal{F}_{c}||\Delta\Theta_{c}|$ and $N_{\text{fine}}=|\mathcal{B}_{f}||\mathcal{F}_{f}||\Delta\Theta_{f}|$ denote the number of candidate parameter tuples evaluated in the coarse and fine search stages, respectively, consistent with the sets defined in Algorithm 1. For each candidate tuple, the dominant computational cost lies in the evaluation of the Fresnel propagation operator $\mathcal{P}_{z}$ (see Eq. (3)). Implementing this operator via the FFT-based Angular Spectrum Method on a transverse aperture grid of size $N_{x}$ (where $N_{x}\gg N$ is chosen to ensure sufficient sampling density and mitigate boundary reflections) yields a propagation complexity of $\mathcal{O}(N_{x}\log N_{x})$. Consequently, the total computational complexity of the proposed algorithm scales as:
@@
It is worth noting that since the traditional focused beam for the Line-of-Sight (LoS) user (UE-2) is fixed during the optimization of the shadowed user (UE-1), its effective channel column $\mathbf{h}_{2}$ can be precomputed once. This reduces the per-iteration cost solely to the update of the Airy beam column $\mathbf{h}_{1}$.
@@
Figure 5: Airy-Null optimization results in the mixed Shadow-Bright scenario. (a) Geometry showing UE-1 in shadow and UE-2 in bright region. (b) Sum-rate vs. steering angle $\Delta\theta$: The peak occurs at $\Delta\theta\approx-2.8^{\circ}$, not at the geometric zero. (c) Condition number $\kappa$ vs. $\theta$: A sharp dip in $\kappa$ coincides with the sum-rate peak, indicating improved orthogonality. (d)-(e) Comparison of Sum-Rate and Condition Number for Trad-All, Airy-Geo, and Airy-Opt strategies. (f) Analog coupling strengths, highlighting the interference suppression capability of the optimized design.